OpenStudy (anonymous):

If y varies inversely as x, and y = 10 as x = 7, find y for the x-value of 10.

OpenStudy (anonymous):

when y varies inversely what do you know about it y = k/x or y =kx?

OpenStudy (anonymous):

I was thinking=k xt being y=kx but im stumped on k? 10=k7 ???

OpenStudy (anonymous):

y =kx is directly

OpenStudy (anonymous):

so you would do the y-k/x? I'm stumped!

OpenStudy (anonymous):

y=k/x *** sry

OpenStudy (anonymous):

ok plug in the x and y then solve for k

OpenStudy (anonymous):

wait nvm. I figured it out... I got y =7

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thanks :)